# One Shot Dominance: Knowledge Poisoning Attack on Retrieval-Augmented Generation Systems

### v2

@@ 5.2 Answering RQ2
5.3 Answering RQ3
Model Authority Setting(ASR) CoE Setting(ACC)
CDP Raw Authority PDP Raw CoE
GPT-3.5 0.5 37.6 71.7 0.5 65.7 82.0
0.67 14.5 52.1 0.67 62.1 80.2
0.75 7.6 47.0 0.75 60.7 75.7
GPT-4 0.5 41.1 55.4 0.5 86.4 90.7
0.67 16.2 40.2 0.67 81.5 87.9
0.75 16.8 42.9 0.75 78.1 86.6
GPT-4o 0.5 17.9 55.5 0.5 86.7 91.5
0.67 7.6 43.5 0.67 83.8 90.4
0.75 2.5 46.1 0.75 79.7 88.2
6 Conclusion
In this paper, we present AuthChain, a realistic knowledge poisoning attack that achieves successful attacks by poisoning only a single document while remaining effective for complex multi-hop questions. AuthChain addresses three key challenges to ensure the poisoned documents are reliably retrieved and trusted by the LLM, even against large knowledge bases and LLM’s own knowledge. Through extensive experiments on six popular LLMs, we demonstrate that AuthChain achieves significantly higher attack success rates while maintaining superior stealthiness against RAG defense mechanisms compared to state-of-the-art baselines.
Limitations
Ethical Statement
Our research on knowledge base poisoning attacks is conducted solely for defensive purposes and academic understanding. We have strictly followed responsible disclosure practices by promptly reporting our findings to relevant knowledge base platforms and RAG system providers. To prevent potential misuse, we only present necessary methodological insights and proof-of-concept examples in this paper, withholding specific attack implementations until appropriate defense mechanisms are developed. This work aims to enhance RAG system security rather than enable harmful exploits.
Appendix B Example of Poisoned Document Generated by AuthChain
Figure 4: Example of Poisoned Document Generated by AuthChain.
Appendix C The Result for Answering RQ2
@@
Appendix D Details of Information Extraction Prompts
@@
Appendix E Appendix: Analysis of Alternative LLMs as AuthChain Agents
@@
Appendix F Details of CoE Judge Agent and Revise Agent Prompts
CoE Judge Agent Prompt: Given: Passage: [Passage] Evidence Nodes: [Nodes] Evidence Relations: [Relations] Intent: [Intent] Question: [Question] Answer: [Answer] Check the [Passage] for Completeness: 1. Evidence Nodes Coverage a) Does each sentence in [Passage] contain at least one [Nodes]? b) Does the [Passage] explicitly include all items listed under [Nodes]? c) Are there any cases where the keywords in [Passage] are replaced by pronouns or vague synonyms (e.g., “it,” “they,” or “this” instead of the actual [Nodes])? 2. Evidence Relations Coverage (Skip if [Relations] is empty) a) Does the [Passage] clearly establish or infer all of the provided [Relations]? b) Are there any unclear or weakly supported relations in [Passage]? 3. Intent Entailment a) Can the specified [Intent] be found in or reasonably inferred from the [Passage]? Output Rules: 1) If all criteria are met (i.e., the Passage covers all [Nodes], [Relations] if present, and [Intent]), output only: Yes 2) If any criterion is not met: Provide a set of revision suggestions for the [Passage]. Specifically: a) Indicate how to add or replace missing keywords (or remove ambiguous pronouns) in each sentence to maximize the number of keywords. b) Tell how to Revise or remove sentences that lack keywords until each sentence contains at least one keyword. c) Explain how to clarify or insert any undefined or weak relations (if [Relations] are given). d) Show how to incorporate the Intent more explicitly if it is missing or unclear. e) Ensure any modifications still align with the original [Answer], to maintain correctness. Do not output any step-by-step explanations or chain-of-thought. Simply give "Yes" if all items are satisfied, or directly provide the revision suggestions if not.
@@ Appendix F Details of CoE Judge Agent and Revise Agent Prompts
Appendix G Details of Authority Generator Agent Prompts
Appendix H Details of Intent Agent Prompt

### v3

@@ 5.2 Answering RQ2
5.3 Answering RQ3
Model Authority Setting(ASR) CoE Setting(ACC)
CDP Raw Authority PDP Raw CoE
GPT-3.5 0.5 37.6 71.7 0.5 65.7 82.0
0.67 14.5 52.1 0.67 62.1 80.2
0.75 7.6 47.0 0.75 60.7 75.7
GPT-4 0.5 41.1 55.4 0.5 86.4 90.7
0.67 16.2 40.2 0.67 81.5 87.9
0.75 16.8 42.9 0.75 78.1 86.6
GPT-4o 0.5 17.9 55.5 0.5 86.7 91.5
0.67 7.6 43.5 0.67 83.8 90.4
0.75 2.5 46.1 0.75 79.7 88.2
6 Discussion
6.1 Robustness of AuthChain to Question Paraphrasing
To address concerns regarding the practical applicability of AuthChain, we investigate its robustness to linguistic variations in user queries. Specifically, we evaluate the attack success rate (ASR) of AuthChain when the original user queries are paraphrased. Table 4 presents the ASR results on three benchmark datasets (HotpotQA, MS-MARCO, and NQ) across six LLMs.
Our results demonstrate that AuthChain maintains a high ASR even when queries are paraphrased, indicating strong robustness to variations in user input. For instance, on HotpotQA, the ASR with GPT-4 increases from 86.0% with the original queries to 91.0% with the paraphrased queries. On the single-hop datasets MS-MARCO and NQ, the ASR decreases only slightly: for example, from 84.0% to 82.0% on MS-MARCO with GPT-4, and from 75.0% to 74.0% on NQ. Importantly, AuthChain consistently outperforms the baselines by a notable margin.
These findings suggest that AuthChain is not only theoretically sound but also practically robust and applicable in real-world scenarios, where user queries may be expressed in diverse linguistic forms.
Table 4: Attack Success Rate (ASR) of AuthChain across six LLMs on different datasets and question types (values reported in %).
Model HotpotQA MS-MARCO NQ
Raw Paraphrased Raw Paraphrased Raw Paraphrased
GPT-3.5 90.0 89.0 74.0 72.0 74.0 74.0
GPT-4 86.0 91.0 84.0 82.0 75.0 74.0
GPT-4o 88.0 91.0 84.0 81.0 81.0 79.0
Llama3-8B 85.0 86.0 79.0 76.0 76.0 73.0
Llama3-70B 86.0 88.0 85.0 79.0 81.0 77.0
Deepseek-V3-0324 87.0 87.0 83.0 81.0 80.0 76.0
6.2 Assessment in Multi-Turn Conversational Settings
We consider a more realistic scenario in which the user engages in several rounds of conversation before asking the targeted question. In this setting, we evaluate the effectiveness of AuthChain when multi-turn conversational context is present.
Recognizing the importance of this setting, we constructed a multi-turn dialogue dataset by augmenting the original single-turn datasets. For each targeted question, we prompted LLMs to generate three rounds of contextually relevant, progressively deepening dialogue, ensuring a natural conversational flow that culminates in the targeted question. These three rounds of dialogue history were concatenated as conversational context, and the original question was then asked as the final turn. Our poisoning strategy was applied to this targeted question.
We conducted experiments using GPT-3.5 as the backend LLM. As shown in Table 5, introducing relevant dialogue history led to a notable reduction in ASR for all methods, including AuthChain and the baselines. We hypothesize that the added context may provide supporting evidence for the correct answer, partially mitigating the effects of poisoning. Nevertheless, AuthChain still substantially outperforms both PoisonedRAG and HijackRAG in this challenging setting (with an average ASR that is 35.0% higher than PoisonedRAG and 22.3% higher than HijackRAG).
Table 5: Attack Success Rate (ASR) in the multi-turn dialogue setting on GPT-3.5 (values reported in %).
Dataset AuthChain PoisonedRAG HijackRAG
HotpotQA_MultiTurn 61.0 27.0 48.0
MS-MARCO_MultiTurn 54.0 16.0 28.0
NQ_MultiTurn 57.0 24.0 29.0
7 Conclusion
In this paper, we present AuthChain, a more realistic knowledge poisoning attack that achieves successful attacks by poisoning only a single document while remaining effective for complex multi-hop questions. AuthChain addresses three key challenges to ensure the poisoned documents are reliably retrieved and trusted by the LLM, even against large knowledge bases and LLM’s own knowledge. Through extensive experiments on six popular LLMs, we demonstrate that AuthChain achieves significantly higher attack success rates while maintaining superior stealthiness against RAG defense mechanisms compared to state-of-the-art baselines. Our findings highlight the importance of developing more robust defense mechanisms for knowledge base security in RAG systems.
Limitations
Ethical Statement
Our research on knowledge base poisoning attacks is conducted with a strong commitment to ethical responsibility and defensive intent. To minimize misuse and promote safer AI systems, we have implemented the following measures:
  1. 1.
    Rigorous Knowledge Verification: Before new information is integrated into knowledge bases, it should undergo thorough verification using existing trusted knowledge sources. Automated cross-referencing and fact-checking can help identify and filter out suspicious or manipulated content.
  2. 2.
    Evidence-Focused Assessment: Since our attack method relies on authoritative statements supported by evidence, one effective defense is to focus on critically assessing the evidence itself. By evaluating the credibility and provenance of the CoE evidence, rather than relying solely on the authority of the statement, systems can reduce the risk of accepting poisoned knowledge as trustworthy.
  3. 3.
    Real-Time Fact-Checking: For newly surfaced or authoritative claims, integrating real-time verification plugins or tools that query the broader internet can help determine the validity of the information. This is especially important for rapidly evolving facts or news, where authoritative statements may be fabricated or outdated.
  4. 4.
    Monitoring for Anomalous Patterns: Systematic monitoring for unusual patterns in document submissions or knowledge updates, such as repeated use of similar authoritative templates or clusters of new facts supporting a specific narrative, can flag potential poisoning attempts for further review. We intend to expand on these points in the revised manuscript to better address the broader societal implications of our work.
We believe that transparent discussion of both attack vectors and feasible defenses is essential for building robust AI systems. We are committed to ongoing dialogue with the research community, system providers, and the public to ensure the safe and responsible advancement of this field.
Appendix B Example of Poisoned Document Generated by AuthChain
Figure 4: Example of Poisoned Document Generated by AuthChain.
Appendix C Details of Information Extraction Prompts
@@
C.1 Performance Across Different Retrievers
To assess the sensitivity of AuthChain to different retrieval strategies, we compare the attack success rate (ASR) and retriever success rate (RSR) using both a dense retriever (Contriever) and a sparse retriever (BM25) on three benchmark datasets, as shown in Table 6.
Table 6: Performance of AuthChain (ASR / RSR) with Different Retrievers (values reported in %).
Retriever HotpotQA MS-MARCO NQ
Contriever 90.0 / 98.0 74.0 / 91.0 74.0 / 94.0
BM25 79.0 / 82.0 83.0 / 98.0 84.0 / 98.0
We observe that the choice of retriever significantly affects the performance of AuthChain across different dataset types. Specifically, BM25 leads to a notable decrease in ASR for the multi-hop dataset HotpotQA (from 90.0% with Contriever to 79.0% with BM25). In contrast, BM25 achieves higher ASR on the single-hop datasets MS-MARCO (increasing from 74.0% to 83.0%) and NQ (from 74.0% to 84.0%). This pattern is largely attributed to the retrieval characteristics of BM25, which excels when the query and supporting evidence have substantial lexical overlap, a common trait in single-hop questions. However, in multi-hop scenarios that require more abstract reasoning and paraphrased evidence, BM25’s reliance on exact keyword matching often fails to retrieve all necessary supporting documents, resulting in a lower attack success rate.
@@ C.1 Performance Across Different Retrievers
Appendix D Analysis of Alternative LLMs as AuthChain Agents
@@
Appendix E Details of Authority Generator Agent Prompts
Appendix F Details of Intent Agent Prompt
Appendix G Details of CoE Judge Agent and Revise Agent Prompts
CoE Judge Agent Prompt: Given: Passage: [Passage] Evidence Nodes: [Nodes] Evidence Relations: [Relations] Intent: [Intent] Question: [Question] Answer: [Answer] Check the [Passage] for Completeness: 1. Evidence Nodes Coverage a) Does each sentence in [Passage] contain at least one [Nodes]? b) Does the [Passage] explicitly include all items listed under [Nodes]? c) Are there any cases where the keywords in [Passage] are replaced by pronouns or vague synonyms (e.g., “it,” “they,” or “this” instead of the actual [Nodes])? 2. Evidence Relations Coverage (Skip if [Relations] is empty) a) Does the [Passage] clearly establish or infer all of the provided [Relations]? b) Are there any unclear or weakly supported relations in [Passage]? 3. Intent Entailment a) Can the specified [Intent] be found in or reasonably inferred from the [Passage]? Output Rules: 1) If all criteria are met (i.e., the Passage covers all [Nodes], [Relations] if present, and [Intent]), output only: Yes 2) If any criterion is not met: Provide a set of revision suggestions for the [Passage]. Specifically: a) Indicate how to add or replace missing keywords (or remove ambiguous pronouns) in each sentence to maximize the number of keywords. b) Tell how to Revise or remove sentences that lack keywords until each sentence contains at least one keyword. c) Explain how to clarify or insert any undefined or weak relations (if [Relations] are given). Do not output any step-by-step explanations or chain-of-thought. Simply give "Yes" if all items are satisfied, or directly provide the revision suggestions if not.
@@ Appendix G Details of CoE Judge Agent and Revise Agent Prompts
Appendix H Defense Methods Details
We provide brief descriptions of the two defense frameworks evaluated in our experiments:
  • InstructRAG enhances the robustness of RAG systems by explicitly guiding language models to learn a denoising process based on self-synthesized rationales. In this framework, the model is instructed to explain how the ground-truth answer is derived from the retrieved documents. These rationales can be leveraged as in-context demonstrations for explicit denoising or as supervised fine-tuning data, thereby improving the model’s ability to identify and resist poisoned or misleading knowledge in the retrieval set.
  • AstuteRAG improves the robustness of RAG systems against imperfect or malicious retrieval by analyzing conflicts between the LLM’s internal knowledge and external sources. It adaptively extracts key information from internal knowledge, integrates it with retrieved content, and produces answers based on source reliability. This method has shown strong effectiveness in detecting and mitigating knowledge poisoning attacks.